Opened 9 years ago

Last modified 6 years ago

#151 assigned flight processing

GB08/08, flight day 127b/2008, Wytham_Woods

Reported by: mark1 Owned by: adbe
Priority: alpha 4 high Milestone: Reprocessing
Component: Processing: general Keywords:
Cc: Other processors:

Description (last modified by mggr)

Data location: ~arsf/arsf_data/in_progress/2008/flight_data/uk/GB08_08-2008_127b_Wytham_Woods

Data arrived from ARSF via initial SATA disk transfer on 12th May 2008

Scientific details: see ~/arsf/arsf_data/in_progress/2008/ARSF_Applications/GB_and_NL/GB08-08 Science Case.pdf

PI: R. Harding

Sensors:

  • Eagle
  • Hawk

Quicklook

TBD.

Change History (14)

comment:1 Changed 9 years ago by mark1

Looking at Application form for GB08/05 suggests that it might be for a different project...doesn't appear to relate to Wytham Woods. Is the project code correct?

comment:2 Changed 9 years ago by mark1

Should probably process upto the applanix stage and see where the GPS data is located (i.e. at Wytham woods or else where). If it is Wytham Woods then maybe contact Andrew Wilson and see if he wants us to process the data or to send him the applanix and raw files.

comment:3 Changed 9 years ago by mark1

At the meeting on 30th May 2008 Gary suggested just sending the Applanix and raw data to Andrew and CC-ing an email to 2 other people at CEH (can't remember who :S)

comment:4 Changed 9 years ago by mark1

Starting to think project code is wrong and should be GB08/08 not GB08/05

comment:5 Changed 9 years ago by mark1

  • Description modified (diff)
  • Summary changed from GB08/05, flight day 127b/2008, Wytham_Woods to GB08/08, flight day 127b/2008, Wytham_Woods

After phone conversation with Gary on 2/6/08 have removed all reference to project GB08/05 and renamed as GB08/08

comment:6 Changed 9 years ago by mark1

Also, all reference to Ross Hill is erroneous and should be R Harding.

comment:7 Changed 9 years ago by mark1

Applanix has been done and is on windows box

comment:8 Changed 9 years ago by mggr

  • Description modified (diff)

Updated PI per #151 comments

comment:9 Changed 9 years ago by mark1

Data has been delivered on the 7th July 2008

comment:10 Changed 8 years ago by mggr

  • Resolution set to fixed
  • Status changed from new to closed

Data sent to NEODC for archiving.

comment:11 Changed 6 years ago by benj

  • Resolution fixed deleted
  • Status changed from closed to reopened

Level 1 data doesn't seem to have gotten from AKW to NEODC. Eagle/Hawk need to be reprocessed and sent to NEODC, please notify Charles George (ctg@…) when done.

comment:12 Changed 6 years ago by benj

  • Owner set to adbe
  • Status changed from reopened to assigned

comment:13 Changed 6 years ago by adbe

Created Delivery

comment:14 Changed 6 years ago by adbe

Ready to deliver at:

/users/rsg/arsf/workspace/GB08_08-2008_127b_Wytham_Woods/processing/delivery/GB08_08-127b-hyperspectral-20120201

Note: See TracTickets for help on using tickets.